OpenStudy (anonymous):

evaluate the expression when p= -24 and q=4. p/2q a) -3 B) 48 C) 3 D) -48

OpenStudy (solomonzelman):

plug in 24 instead of p, and plug in 4 instead of q.

OpenStudy (anonymous):

Your answer is -3

OpenStudy (anonymous):

All you have to do is plug in -24 for P and 4 for Q once that's done divide them.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{p}{2q} }\)\(\Huge\color{blue}{ \sf =\frac{(-24)}{2(4)} }\)\(\Huge\color{blue}{ \sf = \frac{-24}{8}= }\) ( guide the asker please! )

OpenStudy (anonymous):

Which value of x makes this equation true? –8 + x = 0 A. 0 B. –8 C. No value of x will make it true. D. 8

OpenStudy (anonymous):

@SolomonZelman I just did.

OpenStudy (solomonzelman):

only after you handed out the answer.

OpenStudy (anonymous):

8-?=0

OpenStudy (anonymous):

is the same as -8+?=0

OpenStudy (anonymous):

so what is you're answer?

OpenStudy (anonymous):

@SolomonZelman look everyone is here for the same thing maybe you are an stict kind of person but you do your thing and i do mine sounds like a plan? Yeah great! I would appreciate it if you back up off my back about it and i'm saying this in the nicest way possible. You have a great and absolute wonderful day.